he New York Times reported (Feb 17, 1996) that subway ridership declined after a fare increase: “There were nearly four million fewer riders in December 1995, the first full month after the price of token increased 20 cents to $1.60, than in the previous December, a 5.3 percent decline.”

According to the quote above, what is the (approximate) price elasticity of demand in this region of the demand curve?
Answer #1

Price elasticity of demand = Percentage change in demand/Percentage change in price

Percentage change in demand = 5.3%

Percentage change in price = P2-P1/(P2+P1/2)*100 = 1.60-1.40/(1.60+1.40/2)*100 = 13.3%

Price elasticity of demand = 5.3/13.3 = 0.39 or 0.40 approximately

As the elasticty is less than 1 therefore the price elasticity of demand is inelastic.